Maximize Z = 2a + b – 3c + 5d
Subjected to;
a + 7b + 3c + 7d ≤ 46
3a – b + c +2d ≤ 8
2a + 3b – c + d ≤ 10
Where,
a, b, c, and d all are ≥ 0
1
Expert's answer
2020-04-01T16:05:19-0400

Solve this problem by the method of Lagrange multipliers. We have

"\begin{cases}
2a+b-3c+5d\to\max\\
a+7b+3c+7d-46\le 0\\
3a-b+c+2d-8\le 0\\
2a+3b-c+d-10\le 0\\
-a\le 0\\
-b\le 0\\
-c\le 0\\
-d\le 0
\end{cases}"

Then "L(a,b,c,d,\lambda)=\lambda_0(2a+b-3c+5d)+"

"+\lambda_1(a+7b+3c+7d-46)+\lambda_2(3a-b+c+2d-8)+"

"+\lambda_3(2a+3b-c+d-10)-\lambda_4a-\lambda_5b-\lambda_6c-\lambda_7d", where "\lambda=(\lambda_0,\lambda_1,\lambda_2,\lambda_3,\lambda_4,\lambda_5,\lambda_6,\lambda_7)"


If "(a,b,c,d)" is a point of local maximum, we have that there is "\lambda\neq 0" such that

"\lambda_0\ge 0,"

"\lambda_1\le0, \lambda_2\le 0, \lambda_3\le 0, \lambda_4\le 0, \lambda_5\le 0, \lambda_6\le 0, \lambda_7\le 0"

"0=L'_a(a,b,c,d,\lambda)=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4"

"0=L'_b(a,b,c,d,\lambda)=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5"

"0=L'_c(a,b,c,d,\lambda)=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6"

"0=L'_d(a,b,c,d,\lambda)=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7"

"\lambda_1(a+7b+3c+7d-46)=0"

"\lambda_2(3a-b+c+2d-8)=0"

"\lambda_3(2a+3b-c+d-10)=0"

"\lambda_4a=0"

"\lambda_5b=0"

"\lambda_6c=0"

"\lambda_7d=0"

We consider the following cases:

1)"\lambda_0=0"

1.1)"\lambda_1=0"

1.1.1)"\lambda_2=0"

1.1.1.1)"\lambda_3=0"

Then from the equations

"0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4"

"0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5"

"0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6"

"0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7"

we obtain "\lambda=0". This case does not satisfy us.

1.1.1.2)"\lambda_3<0"

From the equation

"0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6"

we obtain that "\lambda_6>0". This case does not satisfy us.

1.1.2)"\lambda_2<0"

From the equations

"0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4"

"0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7"

we obtain

"\lambda_4=3\lambda_2+2\lambda_3<0, \lambda_7=2\lambda_2+\lambda_3<0"

From the equations

"\lambda_4a=0, \lambda_7d=0, \lambda_2(3a-b+c+2d-8)=0"

we obtain

"a=0, d=0, 3a-b+c+2d-8=0" , that is "a=d=0, c=b+8"

From the equations "\lambda_3(3b-c-10)=0, \lambda_5b=0, \lambda_6c=0" we obtain "2\lambda_3(b-9)=0, \lambda_5b=0, \lambda_6(b+8)=0"

Since every system of two equations in "b-9=0, b=0, b+8=0" has no solutions, we obtain that at least two numbers in "\{\lambda_3, \lambda_5, \lambda_6\}" are zeros.

Then the equations

"0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5"

"0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6"

imply "\lambda_2=0" . But it is contradicts with our case "\lambda_2<0".

1.2)"\lambda_1<0"

From the equations

"0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4"

"0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7"

we obtain

"\lambda_4=\lambda_1+3\lambda_2+2\lambda_3<0" and "\lambda_7=7\lambda_1+2\lambda_2+\lambda_3<0"

From the equations

"\lambda_1(a+7b+3c+7d-46)=0, \lambda_4a=0, \lambda_7d=0"

we obtain "a+7b+3c+7d-46=0, a=0, d=0", that is "a=d=0, b=\frac{46-3c}{7}"

From the equations "\lambda_2(3a-b+c+2d-8)=0, \lambda_3(2a+3b-c+d-10)=0,"

"\lambda_5b=0, \lambda_6c=0" we obtain

"\lambda_6c=0, \frac{3\lambda_5}{7}\left(\frac{46}{3}-c\right)=0,"

"\frac{10\lambda_2}{7}\left(c-\frac{51}{5}\right)=0, \frac{16\lambda_3}{7}\left(\frac{17}{4}-c\right)=0"

Since every system of two equations from "c=0, \frac{46}{3}-c=0, c-\frac{51}{5}=0, \frac{17}{4}-c=0" has no solutions, we obtain that at least three numbers from "\{\lambda_6,\lambda_5,\lambda_2,\lambda_3\}" are zeros. Then the equations

"0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5"

"0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6"

imply "\lambda_1=0". It contradicts with our case "\lambda_1<0".

2)"\lambda_0>0". Let "\lambda_0=1".

2.1)"\lambda_1=0"

2.1.1)"\lambda_2=0"

2.1.1.1)"\lambda_3=0"

Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" we obtain "\lambda_4=2\lambda_0>0". It does not satisfy us.

2.1.1.2)"\lambda_3<0"

From "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" and "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "0\ge\lambda_6=-3\lambda_0-\lambda_3=-3-\lambda_3" and "0\ge\lambda_7=5\lambda_0+\lambda_3=5+\lambda_3". System "-3-\lambda_3\le 0", "5+\lambda_3\le 0" does not have solutions.

2.1.2)"\lambda_2<0"

2.1.2.1)"\lambda_3=0"

From "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" we obtain "\lambda_5=\lambda_0-\lambda_2>0". I does not satisfy us.

2.1.2.2)"\lambda_3<0"

2.1.2.2.1)"\lambda_4=0"

From "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" we obtain "0>\lambda_3=-1-1.5\lambda_2" and from "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "0\ge\lambda_7=5+2\lambda_2+\lambda_3=4+0.5\lambda_2". System "4+0.5\lambda_2\le 0, -1-1.5\lambda_2<0" has no solutions.

2.1.2.2.2)"\lambda_4<0"

From "\lambda_2(3a-b+c+2d-8)=\lambda_3(2a+3b-c+d-10)="

"=\lambda_4a=0" we obtain "3a-b+c+2d-8=2a+3b-c+d-10=a=0", that is "a=0, b=\frac{3}{2}(6-d), c=\frac{7}{2}\left(\frac{34}{7}-d\right)"

2.1.2.2.2.1)"d=0". Then "b=9, c=17".

From "\lambda_5b=\lambda_6c=0" we obtain "\lambda_5=\lambda_6=0".

From "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" and "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "\lambda_3=1>0" and "\lambda_2=4>0". It does not satisfy us.

2.1.2.2.2.2)"d=6", then "b=0, c=-4"

From "\lambda_6c=\lambda_7d=0" we obtain "\lambda_6=\lambda_7=0". From "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" and "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_2=-\frac{2}{3}, \lambda_3=-\frac{11}{3}". Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" and "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" we obtain "\lambda_4=-\frac{22}{3}, \lambda_5=-\frac{28}{3}".

So "(0,0,-4,6)" can be a point of global maximum.

2.1.2.2.2.3)"d=\frac{34}{7}", then "b=\frac{12}{7}, c=0"

From "\lambda_5b=\lambda_7d=0" we obtain "\lambda_5=\lambda_7=0". From "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" and "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_2=-2, \lambda_3=-1". Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" and "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "\lambda_4=-6" and "\lambda_6=-4".

So "\left(0,\frac{12}{7},0,\frac{34}{7}\right)" can be a point of global maximum.

2.1.2.2.2.4)"d\not\in\left\{0,6,\frac{34}{7}\right\}" , then "b\neq0, c\neq 0". From "\lambda_5b=\lambda_6c=\lambda_7d=0" we obtain "\lambda_5=\lambda_6=\lambda_7=0". Then from "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5", "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6", "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "1-\lambda_2+3\lambda_3=0", "-3+\lambda_2-\lambda_3=0", "5+2\lambda_2+\lambda_3=0". This sysem has no solutions.

2.2)"\lambda_1<0"

2.2.1)"\lambda_2=0"

2.2.1.1)"\lambda_3=0"

2.2.1.1.1)"\lambda_5=0". From "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" we obtain "\lambda_1=-\frac{1}{7}\lambda_0=-\frac{1}{7}". Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" we obtain "\lambda_4=2\lambda_0+\lambda_1>0". It does not satisfy us.

2.2.1.1.2)"\lambda_6=0". From "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "\lambda_1=\lambda_0>0". It does not satisfy us.

2.2.1.1.3)"\lambda_7=0". From "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_1=-\frac{5}{7}\lambda_0=-\frac{5}{7}". From "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" we obtain "\lambda_4=2\lambda_0+\lambda_1>0". It does not satisfy us.

2.2.1.1.5)"0\not\in\{\lambda_5, \lambda_6, \lambda_7\}". From "\lambda_5b=\lambda_6c=\lambda_7d=0" we obtain "b=c=d=0". Since "\lambda_1(a+7b+3c+7d-46)=0" and "\lambda_1<0", we have "a+7b+3c+7d-46=0", that is "a=46". Since "\lambda_4a=0", we have "\lambda_4=0".

From "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" we obtain "\lambda_1=-2\lambda_0=-2". Then from "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5", "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6", "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_5=-13, \lambda_6=-9, \lambda_7=-9". So "(46,0,0,0)" can be a point of global maximum.

2.2.1.2)"\lambda_3<0"

2.2.1.2.1)"\lambda_5=0". From "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" we obtain "\lambda_3=-\frac{7\lambda_1+1}{3}". Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" we obtain "0\ge\lambda_4=2+\lambda_1+2\lambda_3=\frac{4-11\lambda_1}{3}", that is "\lambda_1\ge\frac{4}{11}>0". It does not satisfy us.

Then "\lambda_5\neq 0". Since, in addition, "\lambda_1\neq 0" and "\lambda_3\neq 0", from "\lambda_1(a+7b+3c+7d-46)="

"=\lambda_3(2a+3b-c+d-10)=\lambda_5b=0" we obtain "a+7b+3c+7d-46="

"=2a+3b-c+d-10=b=0", that is "c=\frac{13}{7}\left(\frac{82}{13}-d\right), a=\frac{10}{7}\left(\frac{38}{5}-d\right)"

2.2.1.2.2)"d=0", then "c=\frac{82}{7}, a=\frac{76}{7}"

Since "\lambda_4a=\lambda_6c=0", we have "\lambda_4=\lambda_6=0". Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" and "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "\lambda_1=\frac{4}{7}>0". It does not satisfy us.

2.2.1.2.3)"d=\frac{82}{13}" , then "c=0, a=\frac{24}{13}"

Since "\lambda_4a=\lambda_7d=0", we have "\lambda_4=\lambda_7=0". Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" and "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_1=-\frac{4}{5}, \lambda_3=-\frac{3}{5}". Then from "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" and "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "\lambda_5=-\frac{32}{5}, \lambda_6=-\frac{24}{5}". So "\left(\frac{24}{13},0,0,\frac{82}{13}\right)" can be a point of global maximum.

2.2.1.2.4)"d=\frac{38}{5}", then "c=-\frac{12}{5}, a=0"

Since "\lambda_6c=\lambda_7d=0", we have "\lambda_6=\lambda_7=0". Then from "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" and "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_3=-\frac{18}{5}, \lambda_1=-\frac{1}{5}".

From "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" and "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" we obtain "\lambda_4=-\frac{27}{5}, \lambda_5=-\frac{56}{5}". So "\left(0,0,-\frac{12}{5},\frac{38}{5}\right)" can be a point of global maximum.

2.2.1.2.5)"d\not\in\left\{0,\frac{82}{13},\frac{38}{5}\right\}", then "a\neq 0, c\neq 0". From "\lambda_4a=\lambda_6c=\lambda_7d=0" we obtain "\lambda_4=\lambda_6=\lambda_7=0". Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4", "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6", "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain that "2+\lambda_1+2\lambda_3=0, -3+3\lambda_1-\lambda_3=0, 0=5+7\lambda_1+\lambda_3". This has no solutions.

2.2.2)"\lambda_2<0"

2.2.2.1)"\lambda_3<0"

From "\lambda_1(a+7b+3c+7d-46)=0", "\lambda_2(3a-b+c+2d-8)=0", "\lambda_3(2a+3b-c+d-10)=0" we obtain "a=-\frac{14}{33}(18+d), b=\frac{29}{66}\left(\frac{94}{29}-d\right), c=\frac{7}{6}\left(\frac{194}{7}-d\right)"

2.2.2.1.1)"d=0", then "a=-\frac{84}{11}, b=\frac{47}{33}, c=\frac{97}{3}"

Since "\lambda_4a=\lambda_5b=\lambda_6c=0", we have "\lambda_4=\lambda_5=\lambda_6=0".

From "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4", "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5", "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "\lambda_1=\frac{16}{33}>0, \lambda_2=\frac{4}{33}>0, \lambda_3=-\frac{47}{33}" . I does not satisfy us.

2.2.2.1.2)"d=-18", then "a=0, b=\frac{28}{3}, c=\frac{160}{3}". Since "\lambda_5b=\lambda_6c=\lambda_7d=0", we have "\lambda_5=\lambda_6=\lambda_7=0".

From "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5", "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6", "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_1=-\frac{2}{7}, \lambda_2=\frac{2}{7}>0, \lambda_3=\frac{3}{7}>0" . I does not satisfy us.

2.2.2.1.3)"d=\frac{94}{29}", then "a=-\frac{784}{27}, b=0, c=\frac{828}{29}"

Since "\lambda_4a=\lambda_6c=\lambda_7d=0", we have "\lambda_4=\lambda_6=\lambda_7=0"

From "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4", "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6", "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_1=-\frac{7}{29}, \lambda_2=\frac{4}{29}>0,\lambda_3=-\frac{63}{58}" . It does not satisfy us.

2.2.2.1.4)"d=\frac{194}{7}", then "a=-\frac{640}{33}, b=-\frac{828}{77}, c=0"

Since "\lambda_4a=\lambda_5b=\lambda_7d=0", we have "\lambda_4=\lambda_5=\lambda_7=0"

From "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4", "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6", "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_1=-\frac{22}{29}, \lambda_2=\frac{54}{29}>0, \lambda_3=-\frac{99}{29}". I does not satisfy us.

2.2.2.1.5)"d\not\in\left\{0,-18,\frac{94}{29},\frac{194}{27}\right\}", then "a\neq 0, b\neq 0, c\neq 0"

Since "\lambda_4a=\lambda_5b=\lambda_6c=\lambda_7d=0", we have "\lambda_4=\lambda_5=\lambda_6=\lambda_7=0"

Then "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4", "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" , "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6", "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7". This system has no solutions.

2.2.2.2)"\lambda_3=0"

2.2.2.2.1)"\lambda_6=0". Then from "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "3\lambda_1+\lambda_2=3". Since "\lambda_1<0, \lambda_2<0" , it does not satisfy us.


Then "\lambda_6<0". From "\lambda_6c=0" we obtain "c=0"

Since "\lambda_1<0, \lambda_2<0", and "\lambda_1(a+7b+3c+7d-46)=0" , "\lambda_2(3a-b+c+2d-8)=0", we obtain "a+7b+3c+7d-46=0, 3a-b+c+2d-8=0"

So "c=0, b=\frac{19}{22}\left(\frac{130}{19}-d\right), a=\frac{21}{22}\left(\frac{34}{7}-d\right)"

2.2.2.2.2)"d=0" , then "b=\frac{65}{11}, a=\frac{51}{11}"

Since "\lambda_4a=\lambda_5b=0", we have "\lambda_4=\lambda_5=0". Since "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" and "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5", we obtain "\lambda_1=-\frac{5}{22}, \lambda_2=-\frac{13}{22}"

Then from "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" we obtain "\lambda_7=\frac{49}{22}>0". It does not satisfy us.

2.2.2.2.3)"d=\frac{130}{19}", then "b=0, a=-\frac{36}{19}"

Since "\lambda_4a=\lambda_7d=0", we have "\lambda_4=\lambda_7=0" . Since "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" and "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7", we obtain "\lambda_1=-\frac{9}{19}, \lambda_2=-\frac{11}{19}"

Then from "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5", "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "\lambda_5=-\frac{33}{19}, \lambda_6=-5". So "\left(-\frac{36}{19},0,0,\frac{130}{19}\right)" can be a point of global maximum.

2.2.2.2.4)"d=\frac{34}{7}" , then "b=\frac{12}{7}, a=0"

Since "\lambda_5b=\lambda_7d=0" , we have "\lambda_5=\lambda_7=0". Since "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5" and "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7", we obtain "\lambda_1=-\frac{1}{3}, \lambda_2=-\frac{4}{3}". Then from "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4" and "0=-3\lambda_0+3\lambda_1+\lambda_2-\lambda_3-\lambda_6" we obtain "\lambda_4=-\frac{7}{3}" and "\lambda_6=-\frac{16}{3}". So "\left(0,\frac{12}{7},0,\frac{34}{7}\right)" can be a point of global maximum.

2.2.2.2.5)"d\not\in\left\{0,\frac{130}{19},\frac{34}{7}\right\}", then "a\neq 0, b\neq 0". Since "\lambda_4a=\lambda_5b=\lambda_7d=0", we have "\lambda_4=\lambda_5=\lambda_7=0". Then the system "0=2\lambda_0+\lambda_1+3\lambda_2+2\lambda_3-\lambda_4", "0=\lambda_0+7\lambda_1-\lambda_2+3\lambda_3-\lambda_5", "0=5\lambda_0+7\lambda_1+2\lambda_2+\lambda_3-\lambda_7" has no solutions.


Consider set of our points "A=\bigl\{(0,0,-4,6),\left(0,\frac{12}{7},0,\frac{34}{7}\right),(46,0,0,0),"

"\left(\frac{24}{13},0,0,\frac{82}{13}\right),\left(0,0,-\frac{12}{5},\frac{38}{5}\right),"

"\left(-\frac{36}{19},0,0,\frac{130}{19}\right),\left(0,\frac{12}{7},0,\frac{34}{7}\right)\bigr\}"

Let "B=\{(a,b,c,d)| a+7b+3c+7d\le 46,"

"3a-b+c+2d \le 8, 2a+3b-c+d\le 10,"

"a\ge 0, b\ge 0, c\ge 0, d\ge 0\}".

Then "A\cap B=\left\{\left(0,\frac{12}{7},0,\frac{34}{7}\right)\right\}"

Let "f(a,b,c,d)=2a+b-3c+5d", then "f(A\cap B)=\{26\}".

We have that "B" is closed as intersection of closed sets. Also "B\subset\{(a,b,c,d)|a+7b+3c+7d\le 46,"

"a\ge 0, b\ge 0, c\ge 0, d\ge 0\}\subset"

"\subset[0,46]\times\left[0,\frac{46}{7}\right]\times\left[0,\frac{46}{3}\right]\times\left[0,\frac{46}{7}\right]". So "B" is bounded. By the Heine–Borel theorem we obtain that "B" is a compact set. Then since "f" is continuous on "B", there is a maximum of "f(B)".

We have that the maximum of "f(B)" is "f\left(0,\frac{12}{7},0,\frac{34}{7}\right)=26".

Answer: "26" in point "\left(0,\frac{12}{7},0,\frac{34}{7}\right)"